OpenStudy (timtim):

V = 100 sin (200πt + π/4) using integral calculus, calculate the RMS value of the voltage

OpenStudy (ash2326):

We know that RMS is root mean square, we'll integrate it over a period Period T=\(\large \frac{2\pi}{200\pi}=\frac{1}{100}\) \[\large V_{rms}=\sqrt{\int_0 ^{1/100} \frac{1}{\frac 1{100}} (100\times \sin (200\pi+\pi/4) )^2dt}\]

OpenStudy (anonymous):

?

OpenStudy (ash2326):

@timtim do you get this?

OpenStudy (anonymous):

with mobile number? ;P

OpenStudy (ash2326):

@timtim now you need it to integrate it, can you do it?

OpenStudy (anonymous):

no he cannt...

OpenStudy (timtim):

not sure how to integrate it why is the period 1/100

OpenStudy (ash2326):

sin t has period of 2pi seconds sin 2t has period of 2pi/2=pi seconds so sin 200\(\pi\)t has period of \(\large \frac{2\pi}{200\pi}=\frac 1{100}\)

OpenStudy (timtim):

ok - understand that bit now so then it needs integrating

OpenStudy (ash2326):

yeah, just integration:)

OpenStudy (ash2326):

brb

OpenStudy (timtim):

10000/100(sin(200pi + pi/4)^2 dt is that right

OpenStudy (ash2326):

yeah

OpenStudy (timtim):

since it is sin^2 do i have to change to cos i.e. 100 x 1/2 [1-cos2(200pi + pi/4)]

OpenStudy (ash2326):

yeah:)

OpenStudy (timtim):

i now have squar route 50[1-cos400pi + pi/2] how do i now integrate

OpenStudy (ash2326):

\[\int \frac{1-\cos (400\pi t+\pi/2)}{2} dt\] \[\frac t2 -\frac{\sin(400\pi t+\pi/2)}{2\times 400 \pi}\]

OpenStudy (ash2326):

apply the limits now

OpenStudy (ash2326):

@timtim do you get this?

OpenStudy (timtim):

substitute t for 1/100 and 0